OpenStudy (anonymous):

A quarterly telephone bill consists of $19.15 rental plus 4.7 cents for each dialled unit. Tax is added at 15%. What is the total bill for Bryndis who used 915 dialled units?

OpenStudy (anonymous):

Convert 4.7 cents to $0.047 amount paid for minutes = 915 * 0.047 Add the rental fee = (915 * 0.047) + 19.15 Tax is 15%, so multiply ^ by 1.15 to get \(total=1.15[(915×0.047)+19.15]\)